OpenStudy (anonymous):

prove the identity from left to right sin^5x = (1-2cos^2x+cos^4x)sinx

OpenStudy (lalaly):

\[\sin^5x= \sin x (\sin ^2 x )^2 \] \[= \sin x(1-\cos^2 x)^2 \] \[= \sin x(1- 2\cos ^2 x + \cos^4 x)\]

OpenStudy (anonymous):

thank you for helpin! you are sooo pretty!

OpenStudy (lalaly):

thats sweet :) Thanks